# Look Twice Before You Answer: Memory-Space Visual Retracing for Hallucination Mitigation in Multimodal Large Language Models

### v1

@@ C.3 Case Study
Figure 11: A case study comparing the levels of hallucination among various baselines
@@
Figure 12: A case study comparing the levels of hallucination among various baselines
@@
Figure 13: A case study comparing the levels of hallucination among various baselines
@@
Figure 14: A case study comparing the levels of hallucination among various baselines
@@
Figure 15: A case study comparing the levels of hallucination among various baselines
@@
Figure 16: A case study comparing the levels of hallucination among various baselines
@@
Figure 17: A case study comparing the levels of hallucination among various baselines
@@
Figure 18: A case study comparing the levels of hallucination among various baselines
@@
Figure 19: A case study comparing the levels of hallucination among various baselines
C.4 Additional Experiments and Results
@@
Method     LLaVABench (in-the-wild)
Average All_1 All_2 All_3
+ OPERA (Huang et al., 2024a) - - - -
+ OPERA (Huang et al., 2024a) - - - -
Table A2: Results on LLaVABench (in-the-wild) dataset. Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline.
Method Total
LLaVA1.5-7B 31.1  $\uparrow$0.0
+ VCD (Leng et al., 2024) 30.20  $\downarrow$0.9
+ OPERA (Huang et al., 2024a) 32  $\uparrow$0.9
+ MemVR (Ours) 32.4  $\uparrow$1.3
Qwen-VL-Chat 49.0  $\uparrow$0.0
+ VCD (Leng et al., 2024) 34.60  $\downarrow$14.4
+ OPERA (Huang et al., 2024a) -
+ MemVR (Ours) 49.6  $\uparrow$0.6
GLM-4V-9B 63.4  $\uparrow$0.0
+ VCD (Leng et al., 2024) 59.40  $\downarrow$4.0
+ OPERA (Huang et al., 2024a) -
+ MemVR (Ours) 65.0  $\uparrow$1.6
Table A3: Results on MM-Vet dataset. Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where lower values indicate better performance.
Method Accuracy
LLaVA1.5-7B 50.00  $\uparrow$0.0
+ VCD (Leng et al., 2024) 44.90  $\downarrow$5.1
+ OPERA (Huang et al., 2024a) 50.76  $\uparrow$0.8
+ MemVR (Ours) 51.50  $\uparrow$1.5
Qwen-VL-Chat 66.05  $\uparrow$0.0
+ VCD (Leng et al., 2024) 34.54  $\downarrow$31.5
+ OPERA (Huang et al., 2024a) -
+ MemVR (Ours) 66.36  $\uparrow$0.3
GLM-4V-9B 57.39  $\uparrow$0.0
+ VCD (Leng et al., 2024) 48.04  $\downarrow$9.4
+ OPERA (Huang et al., 2024a) -
+ MemVR (Ours) 58.00  $\uparrow$0.6
Table A4: Results on Vizwiz dataset. Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where lower values indicate better performance.
Method     CHAIRS
Cs Ci Recall Len
+ OPERA (Huang et al., 2024a) - - - -
+ OPERA (Huang et al., 2024a) - - - -
Method     MMBench-Dev-EN
AR CP FP-C FP-S LR RR Overall
+ VCD (Leng et al., 2024) 60.30 68.58 51.75 53.24 18.64 48.70 54.21
+ OPERA (Huang et al., 2024a) 69.85 75.00 56.64 66.21 28.81 53.04 62.80
+ VCD (Leng et al., 2024) 86.43 85.47 68.53 84.64 61.86 81.74 80.58
+ OPERA (Huang et al., 2024a) - - - - - - -
+ VCD (Leng et al., 2024) 34.67 52.36 20.28 55.63 11.86 22.61 39.18
+ OPERA (Huang et al., 2024a) - - - - - - -
Table A6: Results on MMBench dataset. Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where lower values indicate better performance.
C.5 Supplement Implement Detail
The code of VCD Leng et al. (2024) is also released. However, the result of VCD evaluated in our experiments (e.g. POPE and MME benchmarks) is lower than the original paper. Therefore, we report the results in the original paper.
Appendix D Examples of capability integrations
@@
D.1 Others
Q: How can the model understand information directly from the vision encoder, especially if it has a different vision system?
To ensure that MemVR is adaptable across diverse vision systems, we conducted experiments on multiple VLM architectures, including LLaVA, which utilizes a Visual-Instructional-Tuning framework with different sizes of ViT-based CLIP models, Qwen-VL-Chat, which employs a Q-Former-like architecture for visual processing, and ChatGLM-4v-9B, which integrates a large pre-trained visual encoder. These architectures encompass a broad range of vision models, providing confidence that MemVR is applicable to most VLMs in use today.
Artifacts and licenses
We report a list of licenses for all datasets and models used in our experiment in Table A8. We strictly follow all the model licenses and limit the scope of these models to academic research only.

### v2

@@ B.4 Case Study
Figure 10: A case study comparing the levels of hallucination among various baselines.
@@
Figure 11: A case study comparing the levels of hallucination among various baselines.
@@
Figure 12: A case study comparing the levels of hallucination among various baselines.
@@
Figure 13: A case study comparing the levels of hallucination among various baselines.
@@
Figure 14: A case study comparing the levels of hallucination among various baselines.
@@
Figure 15: A case study comparing the levels of hallucination among various baselines.
@@
Figure 16: A case study comparing the levels of hallucination among various baselines.
@@
Figure 17: A case study comparing the levels of hallucination among various baselines.
@@
Figure 18: A case study comparing the levels of hallucination among various baselines.
B.5 Additional Experiments and Results
Method    MME
Overall Perception Cognition
LLaVA1.5-7B 1864.68  $\uparrow$0.0 1508.97  $\uparrow$0.0 355.71  $\uparrow$0.0
+ VCD (Leng et al., 2024) 1872.87  $\uparrow$8.2 1515.01  $\uparrow$6.0 357.86  $\uparrow$2.2
+ OPERA (Huang et al., 2024a) 1784.34  $\downarrow$80.3 1473.62  $\downarrow$35.3 310.71  $\downarrow$45.0
+ ICD (Wang et al., 2024) 1594.77  $\downarrow$269.9 1306.91  $\downarrow$202.1 287.86  $\downarrow$67.9
+ MemVR (Ours) 1896.72  $\uparrow$32.0 1512.80  $\uparrow$3.8 383.92  $\uparrow$28.2
Qwen-VL-Chat 1784.93  $\uparrow$0.0 1442.79  $\uparrow$0.0 342.14  $\uparrow$0.0
+ VCD (Leng et al., 2024) 1721.06  $\downarrow$63.9 1403.17  $\downarrow$39.6 317.89  $\downarrow$24.3
+ OPERA (Huang et al., 2024a) - - -
+ ICD (Wang et al., 2024) 1833.75  $\uparrow$48.8 1472.06  $\uparrow$29.3 361.69  $\uparrow$19.6
+ MemVR (Ours) 1820.95  $\uparrow$36.0 1473.45  $\uparrow$30.7 347.50  $\uparrow$5.4
GLM-4V-9B 2160.54  $\uparrow$0.0 1680.89  $\uparrow$0.0 479.64  $\uparrow$0.0
+ VCD (Leng et al., 2024) 2105.53  $\downarrow$55.0 1624.10  $\downarrow$56.8 481.43  $\uparrow$1.8
+ OPERA (Huang et al., 2024a) - - -
+ ICD (Wang et al., 2024) 2074.30  $\downarrow$86.2 1566.09  $\downarrow$114.8 508.21  $\uparrow$28.6
+ MemVR (Ours) 2170.17  $\uparrow$9.6 1683.03  $\uparrow$2.1 487.14  $\uparrow$7.5
Table A1: Results on the MME dataset. Best-performing method per model size is in bold. Arrows indicate improvement ( $\uparrow$) or degradation ( $\downarrow$) vs. the baseline model.
Method     LLaVABench (in-the-wild)
Average All_1 All_2 All_3
+ ICD (Wang et al., 2024) 56.90  $\downarrow$7.9 49.80  $\downarrow$13.6 80.70  $\uparrow$0.5 40.20  $\downarrow$10.6
+ OPERA (Huang et al., 2024a) - - - -
+ ICD (Wang et al., 2024) 56.60  $\downarrow$11.9 45.30  $\downarrow$25.1 85.70  $\uparrow$6.4 38.80  $\downarrow$17.0
+ OPERA (Huang et al., 2024a) - - - -
+ ICD (Wang et al., 2024) 70.97  $\downarrow$4.3 82.10  $\downarrow$6.3 71.80  $\downarrow$1.2 59.00  $\downarrow$5.5
Table A2: Results on LLaVABench (in-the-wild) dataset. Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement or degradation over the baseline, where higher values indicate better performance.
Method Total
LLaVA1.5-7B 31.1  $\uparrow$0.0
+ VCD (Leng et al., 2024) 30.2  $\downarrow$0.9
+ OPERA (Huang et al., 2024a) 32.0  $\uparrow$0.9
+ ICD (Wang et al., 2024) 25.9  $\downarrow$5.2
+ MemVR (Ours) 32.4  $\uparrow$1.3
Qwen-VL-Chat 49.0  $\uparrow$0.0
+ VCD (Leng et al., 2024) 34.6  $\downarrow$14.4
+ OPERA (Huang et al., 2024a) -
+ ICD (Wang et al., 2024) 31.7  $\downarrow$17.3
+ MemVR (Ours) 49.6  $\uparrow$0.6
GLM-4V-9B 63.4  $\uparrow$0.0
+ VCD (Leng et al., 2024) 59.4  $\downarrow$4.0
+ OPERA (Huang et al., 2024a) -
+ ICD (Wang et al., 2024) 57.7  $\downarrow$5.7
+ MemVR (Ours) 65.0  $\uparrow$1.6
Table A3: Results on MM-Vet dataset. Best-performing method per model size is highlighted in bold. Arrows indicate improvement ( $\uparrow$) or degradation ( $\downarrow$) over the baseline.
Method Accuracy
LLaVA1.5-7B 50.00  $\uparrow$0.0
+ VCD (Leng et al., 2024) 44.90  $\downarrow$5.1
+ OPERA (Huang et al., 2024a) 50.76  $\uparrow$0.8
+ ICD (Wang et al., 2024) 37.62  $\downarrow$12.4
+ MemVR (Ours) 51.50  $\uparrow$1.5
Qwen-VL-Chat 66.05  $\uparrow$0.0
+ VCD (Leng et al., 2024) 34.54  $\downarrow$31.5
+ OPERA (Huang et al., 2024a) -
+ ICD (Wang et al., 2024) 29.37  $\downarrow$36.7
+ MemVR (Ours) 66.36  $\uparrow$0.3
GLM-4V-9B 57.39  $\uparrow$0.0
+ VCD (Leng et al., 2024) 48.04  $\downarrow$9.4
+ OPERA (Huang et al., 2024a) -
+ ICD (Wang et al., 2024) 50.01  $\downarrow$7.4
+ MemVR (Ours) 58.00  $\uparrow$0.6
Table A4: Results on the Vizwiz dataset. Best-performing method per model size is highlighted in bold; arrows indicate improvement ( $\uparrow$) or degradation ( $\downarrow$) relative to the baseline.
Method     CHAIRS
Cs Ci Recall Len
+ ICD (Wang et al., 2024) 56.20  $\uparrow$8.6 16.30  $\uparrow$3.0 16.30  $\downarrow$64.3 103.40  $\uparrow$3.7
+ OPERA (Huang et al., 2024a) - - - -
+ ICD (Wang et al., 2024) 18.40  $\uparrow$11.6 14.30  $\uparrow$9.0 37.60  $\downarrow$15.8 48.10  $\uparrow$30.5
+ OPERA (Huang et al., 2024a) - - - -
+ ICD (Wang et al., 2024) 43.40  $\uparrow$3.0 9.10  $\uparrow$0.1 73.60  $\uparrow$1.9 239.80  $\uparrow$21.6
Method Existence Count Position Color Scene Artwork OCR Numerical Text_trans Code_reason
LLaVA-Next (Llama3-8B) 195.0 165.0 143.3 185.0 161.6 159.2 118.0 125.0 50.0 77.5
+MemVR 195.0 170.0 143.3 185.0 163.6 161.0 124.0 125.0 52.5 77.5
LLaVA-Next (Mistral-7B) 190.0 150.0 133.3 190.0 144.2 163.5 113.0 122.5 60.0 67.5
+MemVR 195.0 155.0 133.3 190.0 145.2 165.0 113.8 122.5 60.0 67.5
LLaVA-Next (Vicuna-1.6-7B) 195.0 135.0 143.3 165.0 162.2 123.2 132.5 42.5 107.5 55.0
+MemVR 195.0 135.0 135.0 170.0 163.0 123.5 140.0 42.5 115.0 57.5
Table A6: Performance comparison across different LLaVA-Next models with and without MemVR.
Method     MMBench
AR CP FP-C FP-S LR RR Overall
+ VCD (Leng et al., 2024) 60.30 68.58 51.75 53.24 18.64 48.70 54.21
+ OPERA (Huang et al., 2024a) 69.85 75.00 56.64 66.21 28.81 53.04 62.80
+ ICD (Wang et al., 2024) 42.21  $\downarrow$30.7 52.36  $\downarrow$23.3 52.36  $\downarrow$5.7 39.59  $\downarrow$23.9 16.10  $\downarrow$12.7 32.17  $\downarrow$19.1 39.78  $\downarrow$23.0
+ VCD (Leng et al., 2024) 34.67 52.36 20.28 55.63 11.86 22.61 39.18
+ OPERA (Huang et al., 2024a) - - - - - - -
+ ICD (Wang et al., 2024) 12.56  $\downarrow$47.7 17.57  $\downarrow$53.7 2.10  $\downarrow$43.4 22.53  $\downarrow$40.3 2.54  $\downarrow$26.3 5.22  $\downarrow$33.0 13.32  $\downarrow$43.2
+ VCD (Leng et al., 2024) 86.43 85.47 68.53 84.64 61.86 81.74 80.58
+ OPERA (Huang et al., 2024a) - - - - - - -
+ ICD (Wang et al., 2024) 83.92  $\downarrow$4.5 84.46  $\downarrow$2.0 62.94  $\downarrow$7.0 81.23  $\downarrow$4.4 60.17  $\downarrow$5.9 80.00  $\downarrow$5.2 78.01  $\downarrow$4.4
Table A7: Results on the MMBench dataset with newly added ICD rows. Best-performing method per model size and dataset is highlighted in bold; arrows indicate improvement ( $\uparrow$) or degradation ( $\downarrow$) over the baseline.
@@ B.5 Additional Experiments and Results
B.6 VCD Implement Details
The code of VCD (Leng et al., 2024) is also released. However, the result of VCD evaluated in our experiments on POPE is lower than the original paper. Therefore, we report the results in the original paper.
Appendix C Examples of capability integrations
@@
C.1 Others
Q: How can the model understand information directly from the vision encoder, especially if it has a different vision system? To ensure that MemVR is adaptable across diverse vision systems, we conducted experiments on multiple VLM architectures, including LLaVA, which utilizes a Visual-Instructional-Tuning framework with different sizes of ViT-based CLIP models, Qwen-VL-Chat, which employs a Q-Former-like architecture for visual processing, and ChatGLM-4v-9B, which integrates a large pre-trained visual encoder. These architectures encompass a broad range of vision models, providing confidence that MemVR is applicable to most MLLMs in use today.
Artifacts and licenses We report a list of licenses for all datasets and models used in our experiment in Table A10. We strictly follow all the model licenses and limit the scope of these models to academic research only.